Find the following integral. Sx V6x + 7 dx 5x 476x +7 dx = D.
Answer #1

Given: \int x(6x+4)^{\frac{1}{4}}dx

By using Integral by parts :

\int f(x)*g(x)dx=\left [f(x)*\int g(x) \right ]-\int \left [ \frac{\mathrm{d} }{\mathrm{d} x} f(x)*\int g(x)\right ]dx

Here :

  • f(0) = =C
  • g(x)=\left ( 6x+4 \right )^{\frac{1}{4}}

EQUATIONS USED

  • ut ch+1 n+1 +0
  • \int constant*f(x)dx=constant*\int f(x)dx

\Rightarrow \int x(6x+4)^{\frac{1}{4}}dx=\left [ x*\int (6x+4)^{\frac{1}{4}} \right ] -\int \left [ \frac{\mathrm{d} }{\mathrm{d} x}(x)*\int (6x+4)^{\frac{1}{4}} \right ]dx

  • \int (6x+4)^{\frac{1}{4}}dx =\frac{(6x+4)^{\frac{1}{4}+1}}{\frac{1}{4}+1} \div \frac{\mathrm{d} }{\mathrm{d} x}(6x+4)

  =\frac{(6x+4)^{\frac{1+4}{4}}}{\frac{1+4}{4}} \div (6*1+0)

=\frac{(6x+4)^{\frac{5}{4}}}{\frac{5}{4}} \div 6

  =\frac{4}{5}*(6x+4)^{\frac{5}{4}} *\frac{1}{6}

=\frac{2}{5}*(6x+4)^{\frac{5}{4}} *\frac{1}{3}

=\frac{2}{15}(6x+4)^{\frac{5}{4}}

\Rightarrow \int x(6x+4)^{\frac{1}{4}}dx=\left [ x*\frac{2}{15}(6x+4)^{\frac{5}{4}} \right ] -\int \left [1*\frac{2}{15}(6x+4)^{\frac{5}{4}} \right ]dx

  =\left [ \frac{2x}{15}(6x+4)^{\frac{5}{4}} \right ] -\int \left [\frac{2}{15}(6x+4)^{\frac{5}{4}} \right ]dx

  =\left [ \frac{2x}{15}(6x+4)^{\frac{5}{4}} \right ] -\frac{2}{15}\int \left [(6x+4)^{\frac{5}{4}} \right ]dx

  • \int (6x+4)^{\frac{5}{4}}dx =\frac{(6x+4)^{\frac{5}{4}+1}}{\frac{5}{4}+1} \div \frac{\mathrm{d} }{\mathrm{d} x}(6x+4)

  =\frac{(6x+4)^{\frac{5+4}{4}}}{\frac{5+4}{4}} \div (6*1+0)

=\frac{(6x+4)^{\frac{9}{4}}}{\frac{9}{4}} \div 6

  =\frac{4}{9}*(6x+4)^{\frac{9}{4}} *\frac{1}{6}

=\frac{2}{9}*(6x+4)^{\frac{9}{4}} *\frac{1}{3}

=\frac{2}{27}(6x+4)^{\frac{9}{4}}

\Rightarrow \int x(6x+4)^{\frac{1}{4}}dx=\left [ \frac{2x}{15}(6x+4)^{\frac{5}{4}} \right ] -\frac{2}{15}*\frac{2}{27}(6x+4)^{\frac{9}{4}}

\Rightarrow {\color{Red} \int x\sqrt[4]{(6x+4)}dx= \frac{2x}{15}(6x+4)^{\frac{5}{4}} -\frac{4}{405}(6x+4)^{\frac{9}{4}}}